What is the equation for the points (1,5) and (3,6)
Answer by rfer(16322)

m=(6-5)/(3-1)
m=1/2
-----------
y=mx+b
6=.5*3+b
6=1.5+b
b=4.5
----------
y=1/2x+4.5